OpenStudy (fools101):

i need help with one please... Absolute Value Functions Solve: 2|2x + 7| - 3 = 3

OpenStudy (turingtest):

\[2\left| 2x+7 \right|-3=3\]\[2\left| 2x+7 \right|=6\]\[\left| 2x+7 \right|=3\]\[-3=2x+7=3\]\[-10=2x=-4\]\[-5=x=-2\]\[x=\left\{ -2,-5 \right\}\]

OpenStudy (fools101):

okay thank u Turing Test i got that Answer too i was making sure it was right! Thank u!!! : )